OpenStudy (anonymous):

Please help! Simplify: x^2-9/x^2+5x+6

OpenStudy (dumbcow):

Factor: Find factors of -9 that add up to 0 Find factors of 6 that add up to 5

OpenStudy (anonymous):

x+3/x+2?

OpenStudy (dumbcow):

very close x^2-9 = (x+3)(x-3) x^2+5x+6 = (x+3)(x+2) the (x+3) cancels leaving (x-3)/(x+2)

OpenStudy (katrinakaif):

Okay the first one is a perfect square and can be simplified into : (x-3)(x+3) Next simplify the bottom equqation(FACTOR) and get (x+3)(x+2) When you put this togetehr: (x+3)(x-3)/(x+3)(x+2) Cancel out (x+3) = 1 which leaves (x-3)/(x+2)

OpenStudy (anonymous):

Thanks guys! :)